OpenStudy (bishoppatton):

What is 784 x 36 / 14

OpenStudy (pnathaniel56):

answer = 2016

OpenStudy (pnathaniel56):

am i correct

OpenStudy (pnathaniel56):

i get medel

OpenStudy (bishoppatton):

yea... but you used calculator

OpenStudy (pnathaniel56):

no

OpenStudy (pnathaniel56):

give me medele

OpenStudy (bishoppatton):

lies

alones (alones):

@pnathaniel56 you don't force people to give you medals.

alones (alones):

Bishop you need explanation ..how to get 2016?

Awolflover1 (awolflover1):

\[784\times \frac{ 36 }{ 14}=\frac{ 784\times36 }{ 14 }\]

Awolflover1 (awolflover1):

\[784\div14=?\]